# Multi-Agent Inverse Reinforcement Learning in Real World Unstructured Pedestrian Crowds

### v2

@@ Abstract
Social robot navigation in crowded public spaces such as university campuses, restaurants, grocery stores, and hospitals, is an increasingly important area of research. One of the core strategies for achieving this goal is to understand humans’ intent–underlying psychological factors that govern their motion–by learning their reward functions, typically via inverse reinforcement learning (IRL). Despite significant progress in IRL, learning reward functions of multiple agents simultaneously in dense unstructured pedestrian crowds has remained intractable due to the nature of the tightly coupled social interactions that occur in these scenarios e.g. passing, intersections, swerving, weaving, etc. In this paper, we present a new multi-agent maximum entropy inverse reinforcement learning algorithm for real world unstructured pedestrian crowds. Key to our approach is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that achieves tractability at the cost of a slight reduction in accuracy. We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, SCAND, JRDB, and a new dataset, called Speedway, collected at a busy intersection on a University campus focusing on dense, complex agent interactions. Our key findings show that, on the dense Speedway dataset, our approach ranks $1^{\textrm{st}}$ among top $7$ baselines with >2×>2$\times$> 2 × improvement over single-agent IRL, and is competitive with state-of-the-art large transformer-based encoder-decoder models on sparser datasets such as ETH/UCY (ranks $3^{\textrm{rd}}$ among top $7$ baselines).
@@ I Introduction
Robot navigation in densely populated human environments has garnered significant attention, especially in scenarios involving tightly coupled social interactions among pedestrians, scooters, and vehicles. These scenarios [1] occur in both outdoor settings [2, 3, 4, 5] and indoor spaces such as restaurants, grocery stores, hospitals, and university campuses [6, 7, 8, 9, 10, 11]. Effective navigation in such dense and cluttered human environments is challenging requires robots to be socially compliant which involves predicting and responding to human intentions [12, 13]. Many approaches have approached intent prediction implicitly via trajectory forecasting [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. While trajectory forecasting has yielded impressive results on sparse crowd datasets such as ETH, UCY, and JRDB, and structured traffic datasets such as the WAYMO Motion Forecasting Dataset [25], it is unclear whether they perform similarly well in more complex settings that include passing, weaving, yielding to others, etc.
@@
Alternatively, some approaches have modeled intent explicitly by computing humans’ reward functions via inverse reinforcement learning. Recently, Gonon and Billard [26] used maximum entropy IRL [27] for achieving socially compliant navigation in pedestrian crowds. However, inherently a single-agent approach, maximum entropy IRL (MaxEnt IRL) assumes a single objective function for all the agents, which works reasonably well in sparse crowds that lack many tightly coupled interactions, as shown in [26], but fails to accommodate humans’ objectives in denser and more unstructured pedestrian crowds, as we show in our experiments in Section V. This limitation leads us to question: Can MaxEnt IRL be adapted to learn effective robot navigation policies for real-world multi-agent unstructured pedestrian crowds?
Figure 1: Quantitatively benchmarking representative IL approaches on the Speedway dataset: behavior cloning (BC [28]), inverse reinforcement learning (IRL [27] and MAIRL [29]), and generative methods (PECNet [30]). (left) Root mean square error per timestep; larger slope indicates poor long-term planning. (right) Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective learner.
The problem might suggest a multi-agent IRL (Inverse Reinforcement Learning) approach, for which numerous solutions already exist in the dynamic game theory and reinforcement learning literature [31, 32, 33, 34, 35, 36, 37, 38, 29]. The primary limitation to multi-agent IRL is that it assumes known dynamics/system models of all the agents in the environment and the availability of large amounts of training data.
@@
Main contribution: In this work, we propose a novel multi-agent MaxEnt IRL algorithm for learning reward functions of humans in real world unstructured pedestrian crowds. The key insight in our algorithm is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that essentially relaxes the assumption of known system dynamics at the cost of a slight reduction in accuracy. Notably, this trade-off not only makes multi-agent IRL tractable in unstructured settings but also paves the way for new research directions—such as computing exact bounds for this trade-off.
@@
We use off-the-shelf dynamic game theory solvers to roll out trajectories that maximize the learned reward functions. We evaluate this new approach using a dataset collected at a busy intersection on a University campus comprising of pedestrians, scooters, bicycles, particularly focusing on tightly coupled interactions such as when agents walk towards each other, walk past each other, squeeze past each other, etc.We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, JRDB, and a new dataset collected at a busy intersection on a University campus focusing on complex agent interaction. Our findings highlight two key insights: $(i)$ our approach yields the lowest error between the policy-generated trajectories and the ground truth trajectories, $(ii)$ our approach has the highest percentage of low-RMSE trajectories.
@@ II-B Trajectory Prediction
We focus on deep learning-based trajectory and intent prediction which dates back ten years to the seminal SocialLSTM work [42]. TraPHic and RobustTP [43, 44] use an LSTM-CNN framework to predict trajectories in dense and complex environments. TNT [45] uses target prediction, motion estimation, and ranking-based trajectory selection to predict future trajectories. DESIRE [46] uses sample generation and trajectory ranking for trajectory prediction. PRECOG [47] combines conditioned trajectory forecasting with planning objectives for agents. [48] uses a Seq2Seq framework to encode agents’ observations over neighboring agents as their social context for trajectory forecasting and decision-making. [49] uses temporal smoothness in attention modeling for interactions and a sequential model for trajectory prediction. However, a major limitation of this class of methods is that they employ behavior cloning to try and mimic the data, without inherently capturing humans’ internal reward functions.
@@ II-C Intent-aware Multi-agent Reinforcement Learning
Related to our task, intent-aware multi-agent reinforcement learning [50] estimates an intrinsic value that represents opponents’ intentions for communication [51] or decision-making. Many intent inference modules are based on Theory of Mind (ToM) [52] reasoning or social attention-related mechanisms [53, 54]. [55] uses ToM reasoning over opponents’ reward functions from their historical behaviors in performing multi-agent inverse reinforcement learning (MAIRL). [9] uses game theory ideas to reason about other agents’ incentives and help decentralized planning among strategic agents. However, many prior works oversimplify the intent inference and make some prior assumptions about the content of intent.
@@ II-D Opponent Modeling
Opponent modeling [56] has been studied in multi-agent reinforcement learning and usually deploys various inference mechanisms to understand and predict other agents’ policies. Opponent modeling could be done by either estimating others’ actions and safety via Gaussian Process [57] or by generating embeddings representing opponents’ observations and actions [58]. Inferring opponents’ policies helps to interpret peer agents’ actions [59] and makes agents more adaptive when encountering new partners [60]. Notably, many works [61, 62] reveal the phenomenon whereby ego agents’ policies also influence opponents’ policies. To track the dynamic variation of opponents’ strategies made by an ego agent’s influence, [63, 64] propose the latent representation to model opponents’ strategies and influence based on their findings on the underlying structure in agents’ strategy space. [65] provides a causal influence mechanism over opponents’ actions and defines an influential reward over actions with high influence over others’ policies. [66] proposes an optimization objective that accounts for the long-term impact of ego agents’ behavior in opponent modeling. A considerable limitation of many current me thodologies is the underlying assumption that agents continually interact with a consistent set of opponents across episodes.
@@
III Problem Formulation and Background
@@
If $\left(H^{i}_{t}\right)^{-1}$ is non-PSD, perform the Tractability-Rationality Trade-off trick (18)
@@
Figure 2: Flow diagram of our multi-agent inverse reinforcement learning algorithm. The algorithm begins by taking as input the system dynamics, initial state, dataset, feature function, and initial parameters. At each time step, it computes the expected features based on the current policy. These features are used to update the weights by minimizing the difference between the expected and ground truth features. The updated weights are then used to update the game-theoretic policy and roll out a new set of trajectories. If the Hessian matrix is non-positive definite, the ”tractability-rationality trade-off” trick is applied to ensure tractability in unstructured environments.
@@
We define a game, $G\coloneqq\left\langle k,\mathcal{X},\mathcal{T},\{\mathcal{U}^{i}\},\{\mathcal{J}^{i}\}\right\rangle$, where $k$ denotes the number of agents. Hereafter, $i$ will refer to the index of an agent and appear as a superscript whereas $t$ will refer to the current time-step and appear as a subscript. The general state space $\mathcal{X}$ (e.g. SE(2), SE(3), etc.) is continuous; the $i^{\textrm{th}}$ agent at time $t$ has a state $x^{i}_{t}\in\mathcal{X}^{i}$. Over a finite horizon $T$, each agent starts from $x^{i}_{0}\in\mathcal{X}^{i}$ and reaches a goal state $x^{i}_{T}\in\mathcal{X}^{i}$. An agent’s transition function is a mapping, $\mathcal{T}^{i}:\mathcal{X}\times\mathcal{U}^{i}\longrightarrow\mathcal{X}$, where $u^{i}_{t}\in\mathcal{U}^{i}$ is the continuous control space for agent $i$. A discrete trajectory is specified by the sequence $\Gamma^{i}=\left(x^{i}_{0},u^{i}_{0},x^{i}_{1},u^{i}_{1},\ldots,x^{i}_{T-1},u^{i}_{T-1},x^{i}_{T}\right)$.
@@
Each agent has a parameterized running cost [^1] Note that broadly speaking, reward maximization is equivalent to cost minimization, which is what we model in this work. $\mathcal{J}^{i}:\mathcal{X}\times\mathcal{\mathbf{U}}\longrightarrow\mathbb{R}$ where $\mathcal{\mathbf{U}}=\mathcal{U}^{1}\times\mathcal{U}^{2}\times,\ldots,\times\mathcal{U}^{k}$. The cost $\mathcal{J}^{i}$ acts on joint state $x_{t}$ and control $\mathbf{u}_{t}\in\mathbf{U}$ at each time step measuring the distances between agents, deviation from the reference path, and the change in the control input. Each agent $i$ follows a stochastic decentralized policy $\mathcal{F}^{i}_{\theta^{i}}:\mathcal{X}\longrightarrow\mathcal{U}^{i}$, which captures the probability of agent $i$ selecting action $u^{i}_{t}$ at time $t$, given that the system is in state $x_{t}$. The parameters for the policy $\mathcal{F}^{i}$ are represented by $\theta^{i}$, which may correspond to weights of a neural network or coefficients in a weighted linear function. We assume that we are provided with a dataset $\mathcal{\widehat{D}}=\{\widehat{\Gamma}^{1:k}\}_{N}$ that consists of $N$ expert demonstration trajectories, $\widehat{\Gamma}^{i}$ for each agent $i$. Further, we let $\mathcal{D}_{\theta^{i}}=\{\Gamma^{1:k}_{\theta^{i}}\}_{M}$ denote the dataset of $M$ trajectories for each agent generated by the policy $\mathcal{F}^{i}_{\theta^{i}}$ parameterized by $\theta^{i}$ (in the case of IRL, $\mathcal{F}^{i}_{\theta^{i}}$ implies the reward function is parameterized by $\theta^{i}$). Our problem statement is then,
@@
where $\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)$ is a generalized discrepancy function,
$\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)=\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)-\Phi^{i}\left(\widehat{\Gamma}^{i}\right)$ (2)
and $\Phi(\cdot)$ is a feature function that, given a trajectory, returns a vector of features $\left[\phi_{1},\phi_{2},\ldots,\phi_{m}\right]^{\top}$, averaged across the length of the trajectory, where $m$ refers to the number of distinct features. A simple example of a feature function is the weighted linear sum $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)=\sum_{t=0}^{T-1}{\theta^{i}}^{\top}\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$. In IRL, we want to find parameters $\theta^{i}$ for each agent $i$ such that the feature expectation of the generated trajectories $\widehat{\Gamma}^{i}$ under the learned cost weights is as close to the empirical feature expectation of the expert demonstrations. Note that we slightly abuse notation and ignore terminal cost at $t=T$.
III-A Background: MaxEnt IRL
As background, we first briefly review the key details of MaxEnt IRL [26, 27]. Given expert trajectories, the probability of a trajectory $\Gamma^{i}$ under the MaxEnt framework for agent $i$ is defined as:
$P\left(\Gamma^{i}\right)\propto\exp\left(\sum_{t=0}^{T}\mathcal{J}^{i}\left(x_{t},u^{i}_{t}\right)\right)$
where $\mathcal{J}^{i}(x_{t},u^{i}_{t})$ represents the cost for agent $i$ at time $t$ in joint state $x_{t}$ and taking action $u^{i}_{t}$.
The objective in MaxEnt IRL is to adjust the weights of the cost function such that the expected feature counts under the derived policy match the feature counts of the expert, while also maximizing the entropy of the policy. The cost function is iteratively adjusted, typically using gradient-based methods, to ensure that the expected feature counts from the policy induced by the current reward function align with those from the expert demonstrations. Note that since the MaxEnt IRL algorithm traditionally applies to a single agent, Problem (1) is modified to finding $\theta^{*}$ that minimizes
$\left\lVert\mathbb{E}_{\Gamma_{\theta}\sim\mathcal{D}_{\theta}}\Phi\left(\Gamma_{\theta}\right)-\mathbb{E}_{\widehat{\Gamma}\sim\mathcal{\widehat{D}}}\Phi\left(\widehat{\Gamma}\right)\right\rVert$ (3)

### v3

@@ Abstract
Social robot navigation in crowded public spaces such as university campuses, restaurants, grocery stores, and hospitals, is an increasingly important area of research. One of the core strategies for achieving this goal is to understand humans’ intent–underlying psychological factors that govern their motion–by learning how humans assign rewards to their actions, typically via inverse reinforcement learning (IRL). Despite significant progress in IRL, learning reward functions of multiple agents simultaneously in dense unstructured pedestrian crowds has remained intractable due to the nature of the tightly coupled social interactions that occur in these scenarios e.g. passing, intersections, swerving, weaving, etc. In this paper, we present a new multi-agent maximum entropy inverse reinforcement learning algorithm for real world unstructured pedestrian crowds. Key to our approach is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that achieves tractability at the cost of a slight reduction in accuracy. We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, SCAND, JRDB, and a new dataset, called Speedway, collected at a busy intersection on a University campus focusing on dense, complex agent interactions. Our key findings show that, on the dense Speedway dataset, our approach ranks $1^{\textrm{st}}$ among top $7$ baselines with >2×>2$\times$> 2 × improvement over single-agent IRL, and is competitive with state-of-the-art large transformer-based encoder-decoder models on sparser datasets such as ETH/UCY (ranks $3^{\textrm{rd}}$ among top $7$ baselines).
@@ I Introduction
Robot navigation in densely populated human environments has garnered significant attention, especially in scenarios involving tightly coupled social interactions among pedestrians, scooters, and vehicles. These scenarios [1] occur in both outdoor settings [2, 3, 4, 5] and indoor spaces such as restaurants, grocery stores, hospitals, and university campuses [6, 7, 8, 9, 10, 11]. Effective navigation in such dense and cluttered human environments is challenging and requires robots to be socially compliant which involves predicting and responding to human intentions [12, 13]. Many approaches have approached intent prediction implicitly via trajectory forecasting [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. While trajectory forecasting has yielded impressive results on sparse crowd datasets such as ETH, UCY, and JRDB, and structured traffic datasets such as the WAYMO Motion Forecasting Dataset [25], it is unclear whether they perform similarly well in more complex settings that include passing, weaving, yielding to others, etc.
@@
Alternatively, some approaches have modeled intent explicitly by computing humans’ reward functions via inverse reinforcement learning. Recently, Gonon and Billard [26] used maximum entropy IRL (inverse reinforcement learning) [27] for achieving socially compliant navigation in pedestrian crowds. However, inherently a single-agent approach, maximum entropy IRL (MaxEnt IRL) assumes a single objective function for all the agents, which works reasonably well in sparse crowds that lack many tightly coupled interactions, as shown in [26], but fails to accommodate humans’ objectives in denser and more unstructured pedestrian crowds, as we show in our experiments in Section V. This limitation leads us to question: Can MaxEnt IRL be adapted to learn effective robot navigation policies for real-world multi-agent unstructured pedestrian crowds?
We address this question using a multi-agent IRL approach, for which numerous solutions already exist in the dynamic game theory and reinforcement learning literature [28, 29, 30, 31, 32, 33, 34, 35, 36]. Key to making these methods work is the critical assumption of exactly known dynamics/system models of all the agents in the environment, which, however, rarely holds in practice. To relax this assumption, one may consider approximating the dynamics with, say, a simpler model (e.g. constant velocity or constant acceleration models). The downside of such an approximation is that the optimization-based planner does not converge.
@@
Main contribution: In this work, we show that multi-agent MaxEnt IRL provides accurate predictions of pedestrians’ motion in real world unstructured crowds, with only few training trajectories. Our key insight is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that relaxes the assumption of known system dynamics tractably at the cost of a slight reduction in accuracy. Bounded rationality (assumption of agents exhibiting some randomness in their decision-making) is typically, characterized by the covariance matrix of the probabilistic policy. At a high level, the trick uses matrix conditioning (a textbook technique in numerical linear algebra) to improve the condition number of the covariance matrix, thus ensuring tractability, but at the cost of reducing uncertainty, thereby reducing bounded rationality. Notably, this trade-off not only makes multi-agent IRL tractable in unstructured settings but also paves the way for new research directions—such as computing exact bounds for this trade-off.
@@
We use off-the-shelf dynamic game theory solvers to roll out trajectories that maximize the learned reward functions. We evaluate this new approach using a dataset collected at a busy intersection on a University campus comprising of pedestrians, scooters, bicycles, particularly focusing on tightly coupled interactions such as when agents walk towards each other, walk past each other, squeeze past each other, etc. We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, JRDB, and a new dataset collected at a busy intersection on a University campus focusing on complex agent interaction.
In the remainder of this paper, we discuss related work in Section II, formulate the problem in Section III, present our technical approach in Section IV, discuss the experiments and results in Section V, and conclude in Section VI.
@@ II-B Trajectory Prediction
We focus on deep learning-based trajectory and intent prediction which dates back ten years to the seminal SocialLSTM work [40]. TraPHic and RobustTP [41, 42] use an LSTM-CNN framework to predict trajectories in dense and complex environments. TNT [43] uses target prediction, motion estimation, and ranking-based trajectory selection to predict future trajectories. DESIRE [44] uses sample generation and trajectory ranking for trajectory prediction. PRECOG [45] combines conditioned trajectory forecasting with planning objectives for agents. Yoo et al. [46] uses a Seq2Seq framework to encode agents’ observations over neighboring agents as their social context for trajectory forecasting and decision-making. Cao et al. [47] uses temporal smoothness in attention modeling for interactions and a sequential model for trajectory prediction. However, a major limitation of this class of methods is that they employ behavior cloning to try and mimic the data, without inherently capturing humans’ internal reward functions.
@@ II-C Intent-aware Multi-agent Reinforcement Learning
Related to our task, intent-aware multi-agent reinforcement learning [48] estimates an intrinsic value that represents opponents’ intentions for communication [49] or decision-making. Many intent inference modules are based on Theory of Mind (ToM) [50] reasoning or social attention-related mechanisms [51, 52]. Wu et al. [53] uses ToM reasoning over opponents’ reward functions from their historical behaviors in performing multi-agent inverse reinforcement learning (MAIRL). Chandra et al. [9] uses game theory ideas to reason about other agents’ incentives and help decentralized planning among strategic agents. However, many prior works oversimplify the intent inference and make some prior assumptions about the content of intent.
@@ II-D Opponent Modeling
Opponent modeling [54] has been studied in multi-agent reinforcement learning and usually deploys various inference mechanisms to understand and predict other agents’ policies. Opponent modeling could be done by either estimating others’ actions and safety via Gaussian Process [55] or by generating embeddings representing opponents’ observations and actions [56]. Inferring opponents’ policies helps to interpret peer agents’ actions [57] and makes agents more adaptive when encountering new partners [58]. Notably, many works [59, 60] reveal the phenomenon whereby ego agents’ policies also influence opponents’ policies. To track the dynamic variation of opponents’ strategies made by an ego agent’s influence, [61, 62] propose the latent representation to model opponents’ strategies and influence based on their findings on the underlying structure in agents’ strategy space. Jaques et al. [63] provides a causal influence mechanism over opponents’ actions and defines an influential reward over actions with high influence over others’ policies. Kim et al. [64] proposes an optimization objective that accounts for the long-term impact of ego agents’ behavior in opponent modeling. A considerable limitation of many current methodologies is the underlying assumption that agents continually interact with a consistent set of opponents across episodes.
@@
III Background and Problem Formulation
@@
If $\left(H^{i}_{t}\right)^{-1}$ is non-PSD, perform the Tractability-Rationality Trade-off trick (Equation (17))
@@
Figure 1: Flow diagram of the multi-agent inverse reinforcement learning algorithm. The algorithm begins by taking as input the system dynamics, initial state, dataset, feature function, and initial parameters. At each time step, it computes the expected features based on the current policy. These features are used to update the weights by minimizing the difference between the expected and ground truth features. The updated weights are then used to update the game-theoretic policy and roll out a new set of trajectories. If the Hessian matrix is non-positive definite, the “Tractability-Rationality trade-off” trick is applied to ensure tractability in unstructured environments.
@@
We define a game, $G\coloneqq\left\langle k,\mathcal{X},\mathcal{T},\{\mathcal{U}^{i}\},\{\mathcal{J}^{i}\}\right\rangle$, where $k$ denotes the number of agents. Hereafter, $i$ will refer to the index of an agent and appear as a superscript whereas $t$ will refer to the current time-step and appear as a subscript. The general state space $\mathcal{X}$ (e.g. SE(2), SE(3), etc.) is continuous; the $i^{\textrm{th}}$ agent at time $t$ has a state $x^{i}_{t}\in\mathcal{X}^{i}$. Over a finite horizon $T$, each agent starts from $x^{i}_{0}\in\mathcal{X}^{i}$ and reaches a goal state $x^{i}_{T}\in\mathcal{X}^{i}$. An agent’s transition function is a mapping, $\mathcal{T}^{i}:\mathcal{X}\times\mathcal{U}^{i}\longrightarrow\mathcal{X}$, where $u^{i}_{t}\in\mathcal{U}^{i}$ is the continuous control input for agent $i$. A discrete trajectory for agent $i$ is specified by the sequence $\Gamma^{i}=\left(x^{i}_{0},u^{i}_{0},x^{i}_{1},u^{i}_{1},\ldots,x^{i}_{T-1},u^{i}_{T-1},x^{i}_{T}\right)$.
@@
Each agent has a parameterized running cost [^1] Note that broadly speaking, reward maximization is equivalent to cost minimization, which is what we model in this work. $\mathcal{J}^{i}:\mathcal{X}\times\mathcal{{U}}\longrightarrow\mathbb{R}$ where $\mathcal{{U}}=\mathcal{U}^{1}\times\mathcal{U}^{2}\times,\ldots,\times\mathcal{U}^{k}$. The cost $\mathcal{J}^{i}$ acts on joint state $x_{t}$ and control ${u}_{t}\in{U}$ at each time step measuring the distances between agents, deviation from the reference path, and the change in the control input. Each agent $i$ follows a stochastic decentralized policy $\mathcal{F}^{i}_{\theta^{i}}:\mathcal{X}\longrightarrow\mathcal{U}^{i}$, which captures the probability of agent $i$ selecting action $u^{i}_{t}$ at time $t$, given that the system is in state $x_{t}$. The parameters for the policy $\mathcal{F}^{i}$ are represented by $\theta^{i}$, which may correspond to weights of a neural network or coefficients in a weighted linear function. We assume that we are provided with a dataset $\mathcal{\widehat{D}}=\{\widehat{\Gamma}^{1:k}\}$ that consists of expert demonstration trajectories, $\widehat{\Gamma}^{i}$ for each agent $i$. Further, we let $\mathcal{D}_{\theta^{i}}=\{\Gamma^{1:k}_{\theta^{i}}\}$ denote the dataset of trajectories generated by the policy $\mathcal{F}^{i}_{\theta^{i}}$ parameterized by $\theta^{i}$.
III-A Background: MaxEnt IRL
As background, we first briefly review the key details of (single-agent) MaxEnt IRL [26, 27]. As there is only one agent ($k=1$), we omit the superscript $i$ for our discussion on single-agent MaxEnt IRL. Given expert trajectories and assuming that the human is boundedly rational, the probability of a trajectory $\Gamma$ under the MaxEnt IRL framework is defined as:
$P\left(\Gamma\right)\propto\exp\left(\sum_{t=0}^{T}\mathcal{J}\left(x_{t},u_{t}\right)\right)$
where $\mathcal{J}(x_{t},u_{t})$ represents the cost at time $t$ in joint state $x_{t}$ and taking action $u_{t}$.
Let $\Phi(\cdot)$ be a feature function that, given a trajectory, returns a vector of features $\left[\phi_{1},\phi_{2},\ldots,\phi_{m}\right]^{\top}$, averaged across the length of the trajectory, where $m$ refers to the number of distinct features. A simple example of a feature function is the weighted linear sum $\Phi\left(\Gamma_{\theta}\right)=\sum_{t=0}^{T-1}{\theta}^{\top}\mathcal{J}\left(x_{t},{u}_{t}\right)$. The objective in MaxEnt IRL is to adjust the weights of the cost function such that the expected features under the derived policy match the features of the expert, while also maximizing the entropy of the policy. The cost function is iteratively adjusted, typically using gradient-based methods, to ensure that the expected feature counts from the policy induced by the current reward function align with those from the expert demonstrations. To ease notation, let $\Psi\left(\Gamma_{\theta},\widehat{\Gamma}\right)=\Phi\left(\Gamma_{\theta}\right)-\Phi\left(\widehat{\Gamma}\right)$ be a generalized discrepancy function. More formally, MaxEnt IRL finds $\theta^{*}$ that minimizes
$\left\lVert\mathbb{E}_{\widehat{\Gamma},\Gamma_{\theta}\sim\mathcal{\widehat{D}},\mathcal{D}_{\theta}}\Psi\left(\Gamma_{\theta},\widehat{\Gamma}\right)\right\rVert$ (1)
III-B Problem Formulation
We can generalize (1) to the multi-agent setting. Our problem statement is then,
@@ III-B Problem Formulation
Note that we slightly abuse notation and ignore the terminal cost at $t=T$.